x=1,3,4,5,10



P(X=x)= 0.30, 0.20, 0.10, 0.15, 0.25




Question:



a. What is the range of the variable x?



b. What is P(3<X≤10)



c. What is P(X≥3)

1
Expert's answer
2022-03-01T11:00:43-0500

a. Range of the variable x: 10-1=9.


b. "P(3<X\le10)=0.10+0.15+0.25=0.50."


c. "P(X\ge3)=1-0.30=0.70."